OpenStudy (anonymous):

What is the value of m in the following equation: 7 + 2.3m = 20.25 - 3m? 2.5 5.2 2.3 20.25

OpenStudy (texaschic101):

7 + 2.3m = 20.25 - 3 -- subtract 7 from both sides 2.3m = 20.25 - 3 + 7 2.3m = 24.25 -- divide by 2.3 m = 24.25/2.3 m = 10.54 Is the problem written correctly or is there supposed to be an m term on the right side ?

OpenStudy (awesome781):

This is what I was wondering, because I cant get any of the answers listed...

OpenStudy (anonymous):

@texaschic101 I'm sorry, there is supposed to be an m on the right side.

OpenStudy (anonymous):

So it's What is the value of m in the following equation: 7 + 2.3m = 20.25 - 3m?

OpenStudy (texaschic101):

7 + 2.3m = 20.25 - 3m -- subtract 7 from both sides 2.3m = 20.25 - 3m - 7 -- add 3m to both sides 2.3m + 3m = 20.25 - 7 -- combine like terms 5.3m = 13.25 -- divide by 5.3 m = 13.25/5.3 m = 2.5

OpenStudy (anonymous):

Thank you @texaschic101 I like that you included your work so I can learn how to do it :)

OpenStudy (texaschic101):

glad to help :)